A stone is released from rest at a height of 122.5 m. a) What is its velocity after 2 seconds? b) How many seconds will it take for the stone to reach the ground, and c) With what velocity will it hit the ground?

 



1
Expert's answer
2021-09-26T18:47:43-0400

"h=122.5 \;m \\

g = 9.8 \;m/s^2"

a) t= 2 sec

Distance traveled in 2 sec

"S = \frac{1}{2}gt^2 \\

S = \frac{1}{2} \times 9.8 \times 2^2 = 19.6 \;m \\

v^2 -u^2 = 2gS \\

v^2 -0^2 = 2 \times 9.8 \times 19.6 \\

v = \sqrt{384.16} = 19.6 \;m/s"

b) Using the equation of motion

"h=ut+ \frac{1}{2}gt^2 \\

122.5 = 0 \times t + \frac{1}{2} \times 9.8 \times t^2 \\

t^2 = \frac{2 \times 122.5}{9.8} \\

t = \sqrt{25} = 5 \;sec"

c)

"v^2=u^2 +2gh \\

v^2 = 0^2 +2gh \\

v= \sqrt{2gh} \\

v= \sqrt{2 \times 9.8 \times 122.5} \\

v = 49 \;m/s"